If the angles of a triangle are 6x, 4x-20 and 4x-10, what is the measure of each angle of the triangle?
Answer by stanbon(75887)

If the angles of a triangle are 6x, 4x-20 and 4x-10, what is the measure of each angle of the triangle?
----
Equation::
6x + 4x-20 + 4x-10 = 180 degrees
14x -30 = 180
14x = 210
x = 15
------
6x = 6*15 = 90 degrees
4x-20 = 60-20 = 40 degrees
4x-10 = 60-10 = 50 degrees
-----------------
Cheers,
Stan H.
----------